A hypothetical population of 100,000 humans has 68,240 individuals with the blood type AA, 28,735 individuals with blood type AB and-3025 individuals with the blood type BB. a. What is the frequency of each genotype in this population? b. What is the frequency of the A allele? 8. 2p c. What is the frequency B allele? d. If the next generation contained 250,000 individuals, how many individuals would have blood type BB, assuming the population is in Hardy-Weinberg equilibrium?
Answer #1

According to the Hardy-Weinberg equilibrium sum of all the allelic frequency of a gene is always 1 and sum of all the genotypic frequency of all the genotype is always 1.

So p+q =1

p2+ 2pq+q2 =1

here p is the frequency of A allele

q is the frequency of B allele.

A - Frequency of AA genotype = AA population/ total population = 68240 / 100000 = 0.6824

Frequency of AB genotype = AA population/ total population = 28735/ 100000 = 0.28735

Frequency of BB genotype = AA population/ total population = 3025 / 100000 = 0.03025

AA= p2

AB = 2pq

BB = q2

B- Frequency of AA genotype is = 0.6824

Frequency of A allele = V0.6824 = 0.82

C - Frequency of BB genotype is = 0.03025

Frequency of B allele - V0.03025 = 0.18

D- Freqency of BB genotype is 0.03025

So number of individual which has BB genotype will be 250000 * 0.03025 = 7563 individual.